OpenStudy (anonymous):

elimination method for this problem a+8b=17 a+3b=2

OpenStudy (anonymous):

so you would take the a out and do the b 11b=19 right

OpenStudy (ajprincess):

Ya. In this question u can easily eliminate a by subtracting ur second equation from first equation. a+8b=17-(1) a+3b=2-(2) (1)-(2)=> a+8b-a-3b=17-2 5b=15

OpenStudy (ajprincess):

Getting it @bigmommatiad?

OpenStudy (anonymous):

I think I got this one is it (-7, 3)

OpenStudy (ajprincess):

yup u r right:)

OpenStudy (anonymous):

here another one that I need help on 5/4x > -9/18 multiplication of properties inequalities I need to know the solution set or how to do it

OpenStudy (ajprincess):

Is this the question?\[\frac{ 5 }{ 4 }x>\frac{ -9 }{ 18 }\]

OpenStudy (ajprincess):

@bigmommatiad

OpenStudy (anonymous):

sorry, I guess I already did that one, here is myy final question a barrel contains 183 gallon of gas, and it is being drained at a constant rate of 9 gallons a minute. Write equation that models the numberes of gallon, g, after t minutes. The equation that model the given situation is ??